OpenStudy (anonymous):

I have a question about finding the magnitude of a transfer function algebraically.

OpenStudy (anonymous):

The Transfer function is: \[H(s)=(R/L) / (s^2+*(R/L)s+(1/LC)) \] and the equation for the magnitude is: \[|H(jw)|=w(R/L)/(\sqrt{[(1/LC)-w^2]^2+[w(R/L)]^2}\] The problem I am having is remembering how the denominator is manipulated algebraically to separate the complex parts from the real in order to calculate the magnitude.

OpenStudy (anonymous):

sorry the first equation should be... H(s)=(R/L)s/(s2+∗(R/L)s+(1/LC))